OpenStudy (anonymous):

in an examination,65% of total examineesif d total numbr offailures is 420 he total numbr of examinees is

hartnn (hartnn):

65% passed means 100-65 = 35 % failed did u get this?

OpenStudy (anonymous):

how do i solve it

OpenStudy (anonymous):

nut they want d numbr of student

OpenStudy (anonymous):

nut*

OpenStudy (anonymous):

but*

OpenStudy (anonymous):

it wasnt nut it was but

hartnn (hartnn):

in an examination,65% of the total examinees passed if the numbr of failures is 420 the total numbr of examinees is????????????????????

hartnn (hartnn):

so 65% passed how many % failed ?

OpenStudy (anonymous):

yea nd they hav given me options of a)567 b)693 c)1200

OpenStudy (anonymous):

simply u know the % of failures and u want to know 100% then required n of examinees is 100/(100-65) *420 =100/35 *420 =100 *12 =1200

OpenStudy (anonymous):

i dint get it matricked

hartnn (hartnn):

lol @matricked i don't think @aldren understood the word of that

OpenStudy (anonymous):

they dnt want how much % but how many ppl

hartnn (hartnn):

to get how many ppl, we first need to know how much %

OpenStudy (anonymous):

yea exact;y

hartnn (hartnn):

so how much % of ppl failed ?

OpenStudy (anonymous):

35%

OpenStudy (anonymous):

total is 100% now passed is 65% then failed =100-65=35% as per the question 35% is 420 hence 100% is 100/35 * 420 =1200

OpenStudy (anonymous):

thanx alot i understood

OpenStudy (anonymous):

can i ask another 1

hartnn (hartnn):

new question, new post